OpenStudy (anonymous):

Add the complex expressions, and simplify (9 + 9.4i) + (-8.6 – 4i)

OpenStudy (anonymous):

Add real part with real part and imaginary part with imaginary part.

OpenStudy (anonymous):

Are you able to distinguish between real part and imaginary part?

OpenStudy (anonymous):

expand first, 9 + 9.4i -8.6 - 4i then simplify 9.4i - 4i + 9 - 8.6 5.4i + 0.4

OpenStudy (anonymous):

=9-8.6+9.4i-4i =0.4+5.4i

OpenStudy (anonymous):

I am confused...what is real and imaginary? are the ones with the i imaginary?

OpenStudy (anonymous):

Complex numbers are i and real numbers are... well in simple terms, the regular numbers with no i. Where \[i=\sqrt{-1}, i^2=-1, i^3=-i, i^4=1\]

OpenStudy (anonymous):

Ok I got it, thank you! Would you be able to help me with one more?